# The Mass Distribution of Stars Stripped in Binaries: The Effect of Metallicity

### v1

@@
Stars stripped of their hydrogen-rich envelopes through binary interaction are thought to be responsible for both hydrogen-poor supernovae and the hard ionizing radiation observed in low- galaxies. A population of these stars was recently observed for the first time, but their prevalence remains unknown. In preparation for such measurements, we estimate the mass distribution of hot, stripped stars using a population synthesis code that interpolates over detailed single and binary stellar evolution tracks. We predict that for a constant star-formation rate of /yr and regardless of metallicity, a population contains 30,000 stripped stars with mass and 4,000 stripped stars that are sufficiently massive to explode (). Below , the distribution is metallicity-independent and can be described by a power law with the exponent . At higher masses and lower metallicity (), the mass distribution exhibits a drop. This originates from the prediction, frequently seen in evolutionary models, that massive low-metallicity stars do not expand substantially until central helium burning or later and therefore cannot form long-lived stripped stars. With weaker line-driven winds at low metallicity, this suggests that neither binary interaction nor wind mass loss can efficiently strip massive stars at low metallicity. As a result, a “helium-star desert” emerges around at , covering an increasingly large mass range with decreasing metallicity. We note that these high-mass stars are those that potentially boost a galaxy’s He+-ionizing radiation and that participate in the formation of merging black holes. This “helium-star desert” therefore merits further study.
@@
Wolf-Rayet (WR, an observationally-motivated classification) stars are perhaps the most efficient stellar source of ionizing radiation. However, they are predominantly formed through strong stellar winds that are thought to be driven only by metal-rich stars (Crowther, 2007; Smith, 2014). The confirmed existence of WR stars in local metal-poor star-forming galaxies (Massey, 2003) is therefore puzzling, and it has been suggested that these stars are formed instead by envelope-stripping in binaries (Paczyński, 1967; Vanbeveren et al., 1998, see however also Schootemeijer et al. 2024).
@@
While several WR stars have been found orbiting companion stars (Shenar et al., 2016; Shara et al., 2017, 2020), Klencki et al. (2022) showed that massive stars () at low metallicity () may be less likely to create hot, binary-stripped stars. Instead of detaching and contracting to small sizes and high surface temperatures, such stars are predicted to either remain in the mass-transfer phase throughout the rest of their lives (e.g., Bilinski et al., 2024) or only experience partial envelope-stripping, resulting in relatively cool stripped stars (with effective temperatures K). Recent observations support the suggestion that partial stripping could occur (Villaseñor et al., 2023; Ramachandran et al., 2023).
@@
Current stellar evolution models predict that massive stars at low metallicity do not become fully stripped because a large intermediate convection zone develops above the helium core (Schootemeijer et al., 2019; Klencki et al., 2020). The intermediate convection zone efficiently mixes the chemical elements which, combined with the low opacity of stars at low metallicity, allows the envelope to remain small until advanced stages of core-helium burning (see Farrell et al., 2022, for detailed descriptions of the process). When such a star enters into mass transfer, its helium core is already at (or near) thermal equilibrium, which allows for earlier detachment from mass transfer and partial envelope stripping. This late expansion effect is seen in many different stellar evolution models, including MESA (e.g., Schootemeijer et al., 2019), PARSEC (e.g., Tang et al., 2014), GENEC (e.g., Georgy et al., 2013; Groh et al., 2019), and STARS (e.g., Pols et al., 1998), and it becomes increasingly pronounced as metallicity decreases (Klencki et al., 2020). Despite agreement across different models, the late expansion effect is sensitive to uncertain internal mixing processes (Schootemeijer et al., 2019). As a result, this evolutionary phase is an uncertain prediction of stellar models, with potentially large impacts. However, if this late expansion influences the number of massive binary-stripped stars that can form, then observations of the intrinsic rate of such systems could help constrain the internal mixing and radial expansion of massive stars.
@@
The ability (or not) to produce massive binary-stripped stars at low metallicity will impact more than just our understanding of the ionizing radiation from stellar populations. For example, the first evolutionary step in the well-established isolated binary evolution pathway that leads to merging binary black holes and neutron stars as observed by the LIGO-Virgo-Kagra collaboration (Abbott et al., 2023), is successful envelope-stripping through mass transfer in a massive, metal-poor binary star resulting in the creation of a hot, massive helium star (Stevenson et al., 2017; Kruckow et al., 2018). Most rate predictions for merging binary black holes are made with rapid binary population synthesis codes (Mandel & Broekgaarden, 2022), which limits the accuracy with which late-stage mass transfer can be modeled. Furthermore, many of the population models are based on the same single star evolutionary model grid (Pols et al., 1998; Hurley et al., 2000), which necessitates extrapolations for stars more massive than , i.e., the progenitors of most merging black holes observed in gravitational wave merger events. While a detailed modeling study for each separate evolutionary step would be justified, we focus here on the effect of late expansion on the distribution of binary-stripped stars.
@@
In this paper, we model the number and mass distribution of binary-stripped stars as function of metallicity using a binary population synthesis code based on detailed single and binary evolutionary models. In § 2, we describe our binary population synthesis models and our single and binary stellar evolution models that we use in the population models. In § 3, we present our theoretical prediction for the number of stripped stars expected as a function of mass and metallicity—the present-day mass function of stripped stars. § 4 contains estimates for the stripped star populations in the Milky Way, LMC, and SMC, while in § 5, we discuss the implications of our results for high-redshift galaxies and compact object mergers, and we relate our results to other products of massive stars and binaries. Finally, we summarize our conclusions in § 6.
@@ 2 Models
This study is based on binary population synthesis results that rely on binary and single stellar evolution models. In § 2.1, we describe these binary and single stellar evolution models. In § 2.2, we describe the population synthesis code, which simulates a population of stars with properties drawn from observed distributions, then interpolates from the stellar evolution models to determine the properties of the resulting stripped stars.
@@ 2.1 Stellar evolution models
We use the stellar evolution code MESA (version 8118, Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023) to compute two sets of models, one for single stars and one for binaries. These models were first presented in Götberg et al. (2018). For an in-depth discussion of the MESA models, we refer to that manuscript. Here we summarize the most relevant parameters for the current study.
@@ 2.1.1 Evolutionary model inputs
Here, we adopt a common set-up, which instead results in expansion to the red supergiant stage after helium core burning (see for example Sen et al., 2023, but also Georgy et al. 2013; Groh et al. 2019). We assume the Ledoux criterion for semiconvection, setting the efficiency parameter (Langer et al., 1985; Langer, 1991). We implement a step-like convective overshoot, which extends 0.335 pressure scale heights above all convective regions, following the calibrations of Brott et al. (2011) for the extent of the massive star main sequence. To model this in MESA, we use the mixing efficiency from 0.05 pressure scale heights inside the convective region (f = 0.385, f_0 = 0.05).
@@
We adopt standard wind prescriptions during the main-sequence phase (the “Dutch scheme” in MESA). In addition, especially relevant for our models is the wind mass loss during the stripped star phase. For this phase, we extrapolate from the empirical mass loss prescription for WR stars from Nugis & Lamers (2000), again according to the “Dutch scheme” in MESA. Gilkis et al. (2019) studied the effect of adopting the more realistic mass loss rate from Vink (2017) and found that the stripped star could retain more of its hydrogen envelope. When comparing with the results from Gilkis et al. (2019), we find that the stripped star mass could be up to higher than our mass estimates, since we use the stronger mass loss scheme from Nugis & Lamers (2000).
@@
For massive stars with initial masses greater than 20 M⊙, we use the MLT++ mechanism (available in the MESA version 8118 through the control okay_to_reduce_gradT_excess = .true.), an artificial energy transport mechanism used to promote convergence in models with sharp density gradients. MLT++ likely causes our models to underpredict expansion for stars with initial masses (Klencki et al., 2020). However, we expect that the stars with initial masses between 20 and 40 are unaffected. In addition, as described below, the initial masses of the stars in our binary grids do not exceed . As a result, this artificial energy transport only occurs in the single star grids.
@@ 2.1.2 Model grid setup
There are four grids of single star models, each computed at a different metallicity: , , , and . These are designed to resemble stars in the Milky Way, the Large Magellanic Cloud (LMC), the Small Magellanic Cloud (SMC), and the extremely metal-poor star-forming galaxy IZw18, respectively (see, e.g., Szécsi et al., 2015). Their compositions are scaled relative to that of the Sun (Grevesse & Sauval, 1998).
Each grid is made up of 40 models, with initial masses logarithmically spaced between 2 and 100 M⊙. They are computed at least until central helium depletion, thus covering of the stellar lifetime.
To demonstrate late expansion at low metallicity using these single star models, we display in Fig. 1 the radius of single stellar evolution models at the time when helium ignites in the core (see § 2, here we define the helium ignition as when the central helium mass fraction drops below 0.97). While the higher-mass (), high-metallicity () models expand to radii prior to helium ignition, lower metallicity () stars of the same masses do not expand until later, remaining small with radii at this stage. This is similar to the results presented in Klencki et al. (2022).
@@
We model the formation of stripped stars through mass transfer initiated shortly after the donor star finishes the main sequence, early during the Hertzsprung gap. We therefore choose initial periods that range between 3 and 31.7 days and an initial mass ratio of , where is the initial mass of the less massive star in the system. This results in stripped stars with masses () between 0.35 and 8 M⊙, depending on the initial mass and metallicity.
In this study, we use these models to represent all stripped stars that originate from stars with masses between 2 and 18.17 M⊙. For solar metallicity, we do not expect a significant difference in the predicted stripped star mass depending on initial orbital period or initial mass ratio (see e.g., Yungelson et al., 2024). However, at low metallicity, these models could result in a slight underprediction of the stripped star mass (Yoon et al., 2017).
@@ 2.2 Binary stellar population synthesis
We model binary stellar populations using a Monte Carlo code that was developed specifically for studies of stripped stars and was first presented in Götberg et al. (2019). This code uses single and binary stellar evolution models as input, which allows for flexibility in adopting suitable evolutionary models (see also Agrawal et al., 2020, 2023). This method also provides more accuracy in the mass and properties of the stripped stars, compared to the estimate from pure helium star models which are often used in binary population synthesis (Hurley et al., 2000, 2002).
The set-up we use here is similar to what was described in Götberg et al. (2020b). We briefly summarize our assumptions below.
@@
We first draw initial masses () between 0.1 and 100 M⊙, following the initial mass function of Kroupa (2001), in which , with for and for . We determine whether to assign a binary companion to each of these stars following the mass-dependent binary fraction of Moe & Di Stefano (2017, see their Fig. 42). For the binary stars, the initial mass of the less massive star in the system, , is then sampled from a uniform mass ratio distribution, where ranges from 0.1 to 1, in agreement with observed mass ratios (Sana et al., 2012; Kiminki & Kobulnicky, 2012, however, see also Moe & Di Stefano 2017). The initial orbital period () is sampled from a distribution, which is uniform in for binaries with (Öpik, 1924) and favors short initial periods when (, Sana et al., 2012). The minimum initial orbital period sampled is the shortest period for which neither star overflows its Roche lobe at the zero-age main sequence, and the maximum period is days (Moe & Di Stefano, 2017).
@@
We determine when binary interaction is initiated by measuring when the stellar radius, obtained by interpolating MESA models over the initial mass, exceeds the Roche radius (Eggleton, 1983). For simplicity, here we focus on stripped stars produced from the initially most massive star in the system. These should be the most common progenitors for stripped stars, since binaries can be easily disrupted by the creation of the first compact object (e.g., Renzo et al., 2019). We note that stripped stars can be formed from the initially less massive star in a binary system, but they are expected to constitute a smaller fraction (, Zapartas et al., 2017), and since their evolutionary pathways require more detailed scrutiny, we choose to omit their implementation from the population model for this work. Furthermore, we only consider hot stripped stars produced when the Roche lobe is filled prior to core helium burning of the donor star, which we define as when the helium mass fraction in the center first decreases below 0.98. This is a reasonable assumption because envelope-stripping initiated during central helium burning is likely to result in nuclear-timescale mass transfer or cool, partially stripped, puffed-up stars (Klencki et al., 2022, and see § 5.1 for further discussion), and envelope-stripping initiated after central helium depletion will result in very short-lived objects, as only about of the stellar lifetime remains.
@@
For simplicity, we infer the outcome of binary interaction based on limits for the mass ratio and initial period (e.g., POSYDON, Fragos et al. 2023; and BPASS, Eldridge et al. 2017, see also Zapartas et al. 2019). We identify three pathways to the creation of a stripped star, which we depict in Fig. 2. To ensure long-lived stripped stars are created, we only consider donor stars that fill their Roche lobes during either main-sequence evolution (central hydrogen fusion), or during the following rapid expansion phase known as the Hertzsprung gap (hydrogen shell burning after central hydrogen depletion). The three pathways we consider are: (1) stable mass transfer initiated during the main-sequence evolution of the donor star, (2) stable mass transfer initiated during the Hertzsprung gap evolution of the donor star, and (3) the successful ejection of a common envelope initiated during the Hertzsprung gap evolution of the donor star. The Hertzsprung gap typically includes a radial expansion by at least an order of magnitude (see Fig. 1), making it one of the most common evolutionary stages for binary interaction.
We adopt the following approximate approach to determine what systems result in the creation of a stripped star. If the initial mass ratio is and for mass transfer initiated during the main sequence and Hertzsprung gap evolution of the donor star, respectively, we assume that a stripped star can be formed (de Mink et al. 2007; Hurley et al. 2000, see also Zapartas et al. 2019). We assume that common envelopes develop when the donor fills the Roche lobe during the Hertzsprung gap and . Common envelopes that develop earlier in the donor star’s evolution (i.e., on the main sequence) are assumed to lead to coalescence. Common envelopes that develop later in the donor star’s evolution are neglected, meaning that our predicted contribution from the common envelope channel could be conservative. To emulate the result of the common envelope evolution, we adopt the -formalism (Webbink, 1984), setting (e.g., Hurley et al., 2002) and (Dewi & Tauris, 2000; Tauris & Dewi, 2001).
@@
We produce a mapping from initial to stripped star mass from our MESA models and display it in Fig. 3. The different panels show the initial to stripped star mass relation for the different metallicities, as labeled in each top left corner. As shown in the figure, the mapping is somewhat metallicity-dependent. The reason for this metallicity dependence is the lower opacity and weaker winds at low metallicity, which leads to both higher convective core masses during the main sequence and a larger amount of hydrogen left after mass transfer (Yoon et al., 2017; Götberg et al., 2017).
@@
For stars with initial masses less than 20 M⊙, we use our binary-stripped star model grids to determine the mass of the resulting stripped stars (colored, filled circles in Fig. 3 and Sect. 2.1.2). In this case, we adopt the mass of the stripped star halfway through the central helium burning, which we define as when the central helium mass fraction decreases below 0.5. For stars with higher initial masses, we instead use the helium core masses of the single star models at the terminal-age main-sequence (defined as the first time the central hydrogen mass fraction decreases below 0.01), which are marked with colored triangles in Fig. 3 (see Sect. 2.1.2). This estimate gives rise to a maximum stripped star mass of . We choose to use the terminal-age main-sequence core mass because the helium core grows in single stars, while this growth is not prominent for stripped stars. The standard MESA definition for the helium core mass is adopted here, i.e., the location in the stellar interior where the hydrogen mass fraction decreases below 0.1. We note that models of binary-stripped stars typically have more leftover hydrogen on their surfaces, but also that wind mass loss could affect the masses especially at the high-mass end.
@@
In this paper, we do not account for formation of stripped stars through strong stellar winds, i.e., the canonical WR star formation. This suggests that we may underpredict the number of massive helium stars. We expect that wind-stripping would primarily contribute in the high-metallicity regime, where winds are known to be strong (Crowther, 2007).

### v2

@@
11email: beryl@u.northwestern.edu 22institutetext: Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA), Northwestern University, 1800 Sherman Ave, Evanston, IL 60201, USA 33institutetext: Division of Physics, Mathematics, and Astronomy, California Institute of Technology, Pasadena, CA 91125, USA 44institutetext: The Observatories of the Carnegie Institution for Science, 813 Santa Barbara Street, Pasadena, CA 91101, USA 55institutetext: Institute of Science and Technology Austria (ISTA), Am Campus 1, 3400 Klosterneuburg, Austria
55email: ylva.gotberg@ista.ac.at 66institutetext: Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 77institutetext: Max Planck Institute for Astrophysics, Karl-Schwarzschild-Strasse 1, 85748 Garching, Germany 88institutetext: European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85738 Garching bei München, Germany 99institutetext: David A. Dunlap Department of Astronomy and Astrophysics, University of Toronto, 50 St. George Street, Toronto, ON, M5S3H4, Canada
@@
Stars stripped of their hydrogen-rich envelopes through binary interaction are thought to be responsible for both hydrogen-poor supernovae and the hard ionizing radiation observed in low- galaxies. A population of these stars was recently observed for the first time, but their prevalence remains unknown. In preparation for such measurements, we estimate the mass distribution of hot, stripped stars using a population synthesis code that interpolates over detailed single and binary stellar evolution tracks. We predict that for a constant star formation rate of /yr and regardless of metallicity, a scalable model population contains 30,000 stripped stars with mass and 4,000 stripped stars that are sufficiently massive to explode (). Below , the distribution is metallicity-independent and can be described by a power law with the exponent . At higher masses and lower metallicity (), the mass distribution exhibits a drop. This originates from the prediction, frequently seen in evolutionary models, that massive low-metallicity stars do not expand substantially until central helium burning or later and therefore cannot form long-lived stripped stars. With weaker line-driven winds at low metallicity, this suggests that neither binary interaction nor wind mass loss can efficiently strip massive stars at low metallicity. As a result, a “helium-star desert” emerges around at , covering an increasingly large mass range with decreasing metallicity. We note that these high-mass stars are those that potentially boost a galaxy’s He+-ionizing radiation and that participate in the formation of merging black holes. This “helium-star desert” therefore merits further study.
@@
Wolf-Rayet (WR, an observationally motivated classification) stars are perhaps the most efficient stellar source of ionizing radiation. However, in the classical prescription for WR star formation, they are predominantly formed through strong stellar winds that are thought to be driven by metal-rich stars (Crowther, 2007; Smith, 2014, see however Shenar et al. 2020). The confirmed existence of WR stars in local metal-poor star-forming galaxies (Massey, 2003) is therefore puzzling, and it has been suggested that these stars are formed instead by envelope-stripping in binaries (Paczyński, 1967; Vanbeveren et al., 1998, see however also Schootemeijer et al. 2024).
@@
While several WR stars have been found orbiting companion stars (Shenar et al., 2016; Shara et al., 2017, 2020), Klencki et al. (2022) showed that massive stars () at low metallicity () may be less likely to create hot, binary-stripped stars. Instead of detaching and contracting to small sizes and high surface temperatures, such stars are predicted to either remain in the mass-transfer phase throughout the rest of their lives (e.g., Bilinski et al., 2024) or only experience partial envelope-stripping, resulting in relatively cool stripped stars (with effective temperatures K). While these cooler surface temperatures could give them a similar appearance to helium giants (e.g., Laplace et al., 2020), they are expected to be substantially longer lived since they undergo central helium burning at these cooler temperatures. In addition, it is possible that their surface gravities also differ substantially. Recent observations support the suggestion that partial stripping could occur (Villaseñor et al., 2023; Ramachandran et al., 2023).
@@
Current stellar evolution models predict that massive stars at low metallicity do not become fully stripped because a large intermediate convection zone develops above the helium core (Schootemeijer et al., 2019; Klencki et al., 2020). The intermediate convection zone efficiently mixes the chemical elements which, combined with the low opacity of stars at low metallicity, allows the envelope to remain small until advanced stages of core-helium burning (see Farrell et al., 2022, for detailed descriptions of the process). When such a star enters into mass transfer, its helium core is already at (or near) thermal equilibrium, which allows for earlier detachment from mass transfer and partial envelope stripping. This late expansion effect is seen in many different stellar evolution models, including MESA (Modules for Experiments in Stellar Astrophysics, e.g., Schootemeijer et al., 2019), PARSEC (PAdova and TRieste Stellar Evolution Code, e.g., Tang et al., 2014), GENEC (The Geneva Code, e.g., Georgy et al., 2013; Groh et al., 2019), and STARS (e.g., Pols et al., 1998), and it becomes increasingly pronounced as metallicity decreases (Klencki et al., 2020). Despite agreement across different models, the late expansion effect is sensitive to uncertain internal mixing processes (Schootemeijer et al., 2019). As a result, this evolutionary phase is an uncertain prediction of stellar models, with potentially large impacts. However, if this late expansion influences the number of massive binary-stripped stars that can form, then observations of the intrinsic rate of such systems could help constrain the internal mixing and radial expansion of massive stars.
@@
The ability (or not) to produce massive binary-stripped stars at low metallicity will impact more than just our understanding of the ionizing radiation from stellar populations. For example, the first evolutionary step in the well-established isolated binary evolution pathway that leads to merging binary black holes and neutron stars as observed by the LIGO-Virgo-Kagra collaboration (Laser Interferometer Gravitational-Wave Observatory, Abbott et al., 2023), is successful envelope-stripping through mass transfer in a massive, metal-poor binary star resulting in the creation of a hot, massive helium star (Stevenson et al., 2017; Kruckow et al., 2018). Most rate predictions for merging binary black holes are made with rapid binary population synthesis codes (Mandel & Broekgaarden, 2022), which limits the accuracy with which late-stage mass transfer can be modeled. Furthermore, many of the population models are based on the same single star evolutionary model grid (Pols et al., 1998; Hurley et al., 2000), which necessitates extrapolations for stars more massive than , that is, the progenitors of most merging black holes observed in gravitational wave merger events. While a detailed modeling study for each separate evolutionary step would be justified, we focus here on the effect of late expansion on the distribution of binary-stripped stars.
@@
In this paper, we modeled the number and mass distribution of binary-stripped stars as function of metallicity using a binary population synthesis code based on detailed single and binary evolutionary models. In Sect. 2, we describe our binary population synthesis models and our single and binary stellar evolution models that we use in the population models. In Sect. 3, we present our theoretical prediction for the number of stripped stars expected as a function of mass and metallicity—the present-day mass function of stripped stars. Section 4 contains estimates for the stripped star populations in the Milky Way, Large Magellanic Cloud (LMC), and Small Magellanic Cloud (SMC), while in Sect. 5, we discuss the implications of our results for high-redshift galaxies and compact object mergers, and we relate our results to other products of massive stars and binaries. Finally, we summarize our conclusions in Sect. 6.
@@ 2 Models
This study is based on binary population synthesis results that rely on binary and single stellar evolution models. In Sect. 2.1, we describe these binary and single stellar evolution models. In Sect. 2.2, we describe the population synthesis code, which simulates a population of stars with properties drawn from observed distributions, then interpolates from the stellar evolution models to determine the properties of the resulting stripped stars.
@@ 2.1 Stellar evolution models
We used the stellar evolution code MESA (version 8118, Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023) to compute two sets of models, one for single stars and one for binaries. These models were first presented in Götberg et al. (2018). For an in-depth discussion of the MESA models, we refer to that manuscript. Here we summarize the most relevant parameters for the current study.
@@ 2.1.1 Evolutionary model inputs
Here, we adopted a common set-up, which instead results in expansion to the red supergiant stage after helium core burning (see for example Sen et al., 2023, but also Georgy et al. 2013; Groh et al. 2019). We assumed the Ledoux criterion for semiconvection, setting the efficiency parameter (Langer et al., 1985; Langer, 1991). We implemented a step-like convective overshoot, which extends 0.335 pressure scale heights above all convective regions, following the calibrations of Brott et al. (2011) for the extent of the massive star main sequence. To model this in MESA, we used the mixing efficiency from 0.05 pressure scale heights inside the convective region (f = 0.385, f_0 = 0.05).
@@
We adopted standard wind prescriptions during the main-sequence phase (the “Dutch scheme” in MESA). In addition, especially relevant for our models is the wind mass loss during the stripped star phase. For this phase, we extrapolated from the empirical mass loss prescription for WR stars from Nugis & Lamers (2000), again according to the “Dutch scheme” in MESA. Gilkis et al. (2019) studied the effect of adopting the more realistic mass loss rate from Vink (2017) and found that the stripped star could retain more of its hydrogen envelope. When comparing with the results from Gilkis et al. (2019), we found that the stripped star mass could be up to higher than our mass estimates, since we used the stronger mass loss scheme from Nugis & Lamers (2000). Additionally, at stripped star masses , which are contained in our single star model grid but which are more massive than those produced by our binary evolution models, higher mass loss rates than expected from Vink (2017) could be suitable (see Sander & Vink, 2020). We discuss uncertainties that these assumptions could introduce in Sect. 3.1.3.
@@
For massive stars with initial masses greater than 20 M⊙, we used the MLT++ mechanism (available in the MESA version 8118 through the control okay_to_reduce_gradT_excess = .true.), an artificial energy transport mechanism used to promote convergence in models with sharp density gradients. MLT++ likely causes our models to under-predict expansion for stars with initial masses (Klencki et al., 2020). However, we expect that the stars with initial masses between 20 and 40 are unaffected. In addition, as described below, the initial masses of the stars in our binary grids do not exceed . As a result, this artificial energy transport only occurs in the single star grids.
@@ 2.1.2 Model grid setup
There are four grids of single star models, each computed at a different metallicity: , , , and . These are designed to resemble stars in the Milky Way, the LMC, the SMC, and the extremely metal-poor star-forming galaxy IZw18, respectively (see, e.g., Szécsi et al., 2015). Their compositions are scaled relative to that of the Sun (Grevesse & Sauval, 1998).
Each grid is made up of 40 models, with initial masses logarithmically spaced between 2 and 100 M⊙. They were computed at least until central helium depletion, thus covering of the stellar lifetime.
To demonstrate late expansion at low metallicity using these single star models, we display in Fig. 1 the radius of single stellar evolution models at the time when helium ignites in the core (see Sect. 2, here we define the helium ignition as when the central helium mass fraction drops below 0.97). While the higher-mass (), high-metallicity () models expand to radii prior to helium ignition, lower metallicity () stars of the same masses do not expand until later, remaining small with radii at this stage. This is similar to the results presented in Klencki et al. (2022).
At high enough masses, even low metallicity stars are no longer quite so affected by late expansion, and they are able to reach larger radii prior to the onset of core helium burning. Additionally, at these high masses, the higher metallicity stars have very strong winds, causing their evolutionary track in the Hertzsprung Russell diagram to turn around before they reach the Hayashi track, on the way to becoming WR stars. This is why the highest mass stars in Fig. 1 have similar radii regardless of metallicity.
We note that the number of stripped stars estimated in this work is dependent on the radius evolution of the underlying evolutionary model. The radius evolution for high mass stars () can be affected by luminous blue variable-like eruptions and the Humphreys-Davidson limit (e.g., Gilkis et al., 2021; Cheng et al., 2024), which we did not account for in this work.
@@
For our binary star grids, we modeled the formation of stripped stars through mass transfer initiated shortly after the donor star finishes the main sequence, early during the Hertzsprung gap. We therefore chose initial periods that range between 3 and 31.7 days and an initial mass ratio of , where is the initial mass of the less massive star in the system. This resulted in stripped stars with masses () between 0.35 and 8 M⊙, depending on the initial mass and metallicity. Although the MESA binary models are limited to one formation channel of stripped stars, we used them to approximate the result of envelope-stripping through other evolutionary channels as well (see Sect. 2.2).
In this study, we used these models to represent all stripped stars that originate from stars with masses between 2 and 18.17 M⊙. For solar metallicity, we do not expect a significant difference in the predicted stripped star mass depending on initial orbital period or initial mass ratio (see e.g., Yungelson et al., 2024). However, at low metallicity, these models could result in a slight under-prediction of the stripped star mass (Yoon et al., 2017).
@@ 2.2 Binary stellar population synthesis
We modeled binary stellar populations using a Monte Carlo code that was developed specifically for studies of stripped stars and was first presented in Götberg et al. (2019). This code uses single and binary stellar evolution models as input, which allows for flexibility in adopting suitable evolutionary models (see also Agrawal et al., 2020, 2023). This method also provides more accuracy in the mass and properties of the stripped stars, compared to the estimate from pure helium star models which are often used in binary population synthesis (Hurley et al., 2000, 2002).
The set-up we used here is similar to what was described in Götberg et al. (2020b). We briefly summarize our assumptions below.
@@
We first drew initial masses () between 0.1 and 100 M⊙, following the initial mass function of Kroupa (2001), in which , with for and for . We determined whether to assign a binary companion to each of these stars following the mass-dependent binary fraction of Moe & Di Stefano (2017, see their Fig. 42). For the binary stars, the initial mass of the less massive star in the system, , was then sampled from a uniform mass ratio distribution, where ranges from 0.1 to 1, in agreement with observed mass ratios (Sana et al., 2012; Kiminki & Kobulnicky, 2012, however, see also Moe & Di Stefano 2017). The initial orbital period () was sampled from a distribution which is uniform in for binaries with (Öpik, 1924) and favors short initial periods when (, Sana et al., 2012). The minimum initial orbital period sampled is the shortest period for which neither star overflows its Roche lobe at the zero-age main sequence, and the maximum period is days (Moe & Di Stefano, 2017).
@@
We determined when binary interaction is initiated by measuring when the stellar radius, obtained by interpolating MESA models over the initial mass, exceeds the Roche radius (Eggleton, 1983). For simplicity, here we focused on stripped stars produced from the initially most massive star in the system. These should be the most common progenitors for stripped stars, since binaries can be easily disrupted by the creation of the first compact object (e.g., Renzo et al., 2019). We note that stripped stars can be formed from the initially less massive star in a binary system, but they are expected to constitute a smaller fraction (, Zapartas et al., 2017), and since their evolutionary pathways require more detailed scrutiny, we chose to omit their implementation from the population model for this work. Furthermore, we only considered hot stripped stars produced when the Roche lobe is filled prior to core helium burning of the donor star, which we define as when the helium mass fraction in the center first decreases below 0.98. This is a reasonable assumption because envelope-stripping initiated during central helium burning is likely to result in nuclear-timescale mass transfer or cool, partially stripped, puffed-up stars (Klencki et al., 2022, and see Sect. 5.1 for further discussion), and envelope-stripping initiated after central helium depletion will result in very short-lived objects, as only about of the stellar lifetime remains.
@@
For simplicity, we inferred the outcome of binary interaction based on limits for the mass ratio and initial period (e.g., POSYDON, POpulation SYnthesis with Detailed binary-evolution simulatiONs, Fragos et al. 2023; and BPASS, Binary Population and Spectral Synthesis, Eldridge et al. 2017, see also Zapartas et al. 2019). We identified three pathways to the creation of a stripped star, which we depict in Fig. 2. To ensure long-lived stripped stars are created, we only considered donor stars that fill their Roche lobes during either main-sequence evolution (central hydrogen fusion), or during the following rapid expansion phase known as the Hertzsprung gap (hydrogen shell burning after central hydrogen depletion). The three pathways we considered are: (1) stable mass transfer initiated during the main-sequence evolution of the donor star (also known as Case A), (2) stable mass transfer initiated during the Hertzsprung gap evolution of the donor star (Case B), and (3) the successful ejection of a common envelope initiated during the Hertzsprung gap evolution of the donor star (Case B CEE). The Hertzsprung gap typically includes a radial expansion by at least an order of magnitude (see Fig. 1), making it one of the most common evolutionary stages for binary interaction.
We adopted the following approximate approach to determine what systems result in the creation of a stripped star. If the initial mass ratio is and for mass transfer initiated during the main sequence and Hertzsprung gap evolution of the donor star, respectively, we assumed that a stripped star can be formed (de Mink et al. 2007; Hurley et al. 2002, see also Zapartas et al. 2019). We assumed that common envelopes develop when the donor fills the Roche lobe during the Hertzsprung gap and (Hurley et al., 2002). Common envelopes that develop earlier in the donor star’s evolution (i.e., on the main sequence) were assumed to lead to coalescence. Common envelopes that develop later in the donor star’s evolution were neglected, meaning that our predicted contribution from the common envelope channel could be conservative. To emulate the result of the common envelope evolution, we adopted the -formalism (Webbink, 1984), setting (e.g., Hurley et al., 2002) and (Dewi & Tauris, 2000; Tauris & Dewi, 2001). We note that our treatment for common envelope ejection is simplified, since the binding energy of a stellar envelope evolves during expansion (see e.g., Klencki et al., 2021). This means that our predicted contribution from common envelope ejections is somewhat uncertain (see Sect. 3.1.3).
@@
We produced a mapping from initial to stripped star mass from our MESA models and display it in Fig. 3. The different panels show the initial to stripped star mass relation for the different metallicities, as labeled in each top left corner. As shown in the figure, the mapping is somewhat metallicity-dependent. The reason for this metallicity dependence is the lower opacity and weaker winds at low metallicity, which leads to both higher convective core masses during the main sequence and a larger amount of hydrogen left after mass transfer (Yoon et al., 2017; Götberg et al., 2017).
@@
For stars with initial masses less than 20 M⊙, we used our binary-stripped star model grids to determine the mass of the resulting stripped stars (colored, filled circles in Fig. 3 and Sect. 2.1.2). In this case, we adopted the mass of the stripped star halfway through the central helium burning, which we defined as when the central helium mass fraction decreases below 0.5. For stars with higher initial masses, we instead used the helium core masses of the single star models at the terminal-age main-sequence (defined as the first time the central hydrogen mass fraction decreases below 0.01), which are marked with colored triangles in Fig. 3 (see Sect. 2.1.2). This estimate gives rise to a maximum stripped star mass of . We chose to use the terminal-age main-sequence core mass because the helium core grows in single stars, while this growth is not prominent for stripped stars. The standard MESA definition for the helium core mass was adopted here, that is, the location in the stellar interior where the hydrogen mass fraction decreases below 0.1. We note that models of binary-stripped stars typically have more leftover hydrogen on their surfaces, but also that wind mass loss could affect the masses especially at the high-mass end.
@@
In this paper, we did not account for formation of stripped stars through strong stellar winds, that is, the canonical WR star formation. This suggests that we may under-predict the number of massive helium stars. We expect that wind-stripping would primarily contribute in the high-metallicity regime, where winds are known to be strong (Crowther, 2007).